Highest Energy Data

Typically look at energy distribution of “jets” produced in these collisions

• Compare with predictions that assume point quarks

High Energy Data

The high E_{T} tail does show an excess

• Best seen when comparing relative difference

One sees a small discrepancy

• Looks suspicious, but not conclusive

• Implies quark is smaller than ∼ 10−17 m

Why MUST the Top Quark Exist?

If top doesn’t exist, the Standard Model crumples

• Masses of the W ± and Z∘ bosons cannot be explained

– Measurements of the Z boson require

Mtop = 166{+17+19 \atop −19−22} GeV/c2

• Behaviour of the bottom quark implies that it must have a partner

– Weak force predicts asymmetry in e+ e− → bb̄

• B meson “mixing” could not occur

Why are Heavy Quarks Hard to Find?

Hard to produce

• Must sift through lots of matter-antimatter collisions

• Happens about 1 in every billion collisions?

They Decay!

• A top quark decays via the weak force into a bottom quark • The bottom quark then decays into lighter quarks

Ya can’t see free quarks

• Quarks are “not allowed” to appear unescorted in public • Have to “dress” themselves up around other quarks

• Result in “jets” of particles

Now Have the Evidence...

In about 10 billion collisions, find about 50 events

We observe about the right number of different types

• Observe 1 μ+ μ− event and 5 eμ events

• Expect 1.3 ± 0.3 events from background

• Observe ∼ 40 events with W + ≥ 3 jets

Probability of background looking like this is < 10^{−6}